# Conduct Parameter Estimation in Homogeneous Goods Markets with Equilibrium Existence and Uniqueness Conditions: The Case of Log-linear Specification

### v1

@@ Acknowledgments
We thank Jeremy Fox, Yelda Gungor, and Isabelle Perrigne for valuable comments. This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
@@ Appendix A Online appendix
A.1 Existence and uniqueness of equilibrium prices
@@
An important note is that $P_{t}=0$ satisfies $\Delta(P_{t})=0$. Thus $P_{t}^{*}=0$ holds. However, our interest is a unique positive equilibrium price, so we seek conditions in which there is a unique positive price that satisfies $\Delta(P_{t})>0$.
When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, (I) is always negative on $P_{t}>0$. In contrast, (II) is non-negative regardless of the sign of $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ on $P_{t}>0$. Therefore, $\Delta(P_{t})$ is always negative on $P_{t}>0$, which implies that there is no positive equilibrium price.
@@
Therefore, the shape of (II) on $P_{t}>0$ is changed based on the value of $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ and the number that (I) and (II) cross on $P_{t}>0$ is also determined.
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})<0$ is illustrated in Figure 1(a). In this case, (II) is a monotone decreasing convex function because the first and second derivatives are both negative. As (I) is monotone increasing in $P_{t}>0$, (I) and (II) cross only once on $P_{t}>0$.
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\in[0,1)$ is illustrated in Figure 1(b). In this case, (II) becomes a monotone increasing concave function passing through the origin because the first derivative is positive and the second derivative is negative. As the value of the first derivative is infinity at $P_{t}=0$ since $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})-1<0$, (II) should be greater than (I) around $P_{t}=0$. Then as $P_{t}$ becomes large starting from zero, the difference in both terms becomes small, and both terms cross only once on $P_{t}>0$ eventually.
When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, (16) becomes
@@
If $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})=\exp(\Xi)$, the above equation holds for any $P_{t}$. Thus there are infinitely many equilibrium prices on $P_{t}>0$, which is illustrated in Figure 1(c). When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq\exp(\Xi)$, there is no equilibrium price on $P_{t}>0$ because the right-hand side of the equation is always non-zero on $P_{t}>0$, which is illustrated in Figure 1(d).
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})>1$ is illustrated in Figure 1(e). In this case, (II) becomes a monotone increasing convex function passing through the origin because the first and second derivatives are positive. Therefore, (I) and (II) cross only once on $P_{t}>0$.
@@
$P_{t}$$\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$$[1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})]P_{t}$
@@
Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. The thick line represents the first term and the dotted line represents the second term in (16).
@@
00 0 $=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})]P_{t}-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$
00 0 $=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$
@@
Figure 2 illustrates how the demand and supply equations cross under the conditions. When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the supply equation (5) is ill-defined because the inside of the log function becomes negative. Thus, there should not be any equilibrium. Hereafter, assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand equation and supply equation coincides. Thus there are infinitely many equilibria in the model. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand and supply equations have a same slope and different intercepts, which means that both equations become parallel. Thus, there is no equilibrium. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, the demand and supply equations have different slopes, and hence we can find a unique equilibrium.
@@
[…]
(a) $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
[…]
(b) $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
[…]
@@
Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. Case (a) and (b) also allow both the demand and supply equations to be downward sloping.
A.2 Simulation and estimation procedure
@@
We compute the equilibrium quantity $Q_{t}$ for the log-linear model by (20). We then compute the equilibrium price $P_{t}$ by substituting $Q_{t}$ and other variables into the demand function (2). We generate 1000 data sets of 100, 200, 1000, 1500 markets. We jointly estimate the demand and supply parameters by the simultaneous equation model (Wooldridge 2010) from the true values. We use state-of-the-art constrained optimization solvers, i.e., Ipopt.jl which implements an interior point line search filter method that aims to find a local solution of nonlinear programming problems.
@@
A.3 Why does the estimation without equilibrium conditions converge to the extremely low conduct parameters?
We should interpret reasons why the estimation without equilibrium conditions converges to the negative conduct parameters, as shown in Panel (a) in Table 1, although the obtained parameters cannot be interpreted as the exogenous elements determining equilibrium outcomes. First, we rewrite the terms in Equation (7) into
$\log(1-\theta C_{t})-\gamma_{0}$
where $C_{t}=\alpha_{1}+\alpha_{2}Z^{R}_{t}$ is given in the demand estimation. If the model allows $\theta<0$, then $1-\theta C_{t}$ can take an arbitrarily large value through log-transformation relative to $\gamma_{0}$. For example, if $0001-\theta C_{t}=1,000,$ 000, then $\log(1-\theta C_{t})=13.8$, which has a small contribution relative to $\gamma_{0}$ for evaluating the moment conditions.
A.4 Additional experiments
@@
Additional results for different $\sigma$ are shown in Tables 3 and 4. As a summary, the main findings in the main text are robust.
@@
In Tables 6 and 7, we illustrate that N2SLS with Constraints (12) works for the linear model as in Matsumura and Otani (2023b). This means that incorporating the conditions is innocuous for the linear model.
Additional ad hoc improvement is possible by using (3) to compute $\varepsilon_{t}^{c}$ and (4) as constraints with Constraints (12), (13), and (14). Table 5 shows that the estimation of the conduct parameter $\theta$ improves, that is, the bias is 0.014 and RMSE is 0.217. However, the results with ad hoc improvement do not dominate the results without the improvement for all parameters.
@@
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$ 0.119 4.055 0.346 3.180 0.344 1.497 0.379 1.582
$\alpha_{1}$ 0.019 0.620 0.054 0.486 0.076 0.429 0.099 0.576
$\alpha_{2}$ 0.000 0.050 0.003 0.039 0.002 0.029 0.003 0.036
$\alpha_{3}$ 0.009 0.336 0.025 0.268 0.028 0.129 0.031 0.129
$\gamma_{0}$ 10.104 11.066 9.397 10.288 13.240 13.280 13.314 13.346
$\gamma_{1}$0.146 0.340 -0.166 0.264 -0.180 0.198 -0.181 0.192
$\gamma_{2}$0.072 0.229 -0.079 0.165 -0.087 0.108 -0.086 0.099
$\gamma_{3}$0.069 0.230 -0.082 0.171 -0.088 0.110 -0.087 0.100
$\theta$5e+04 1e+05 -3e+04 6e+04 -8e+04 9e+04 -8e+04 9e+04
Runs converged (%) 100.000 100.000 98.100 98.700
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.006 3.556 0.175 3.158 0.059 0.977 -0.017 0.763
$\alpha_{1}$0.002 0.545 0.026 0.484 0.009 0.150 -0.003 0.117
$\alpha_{2}$ 0.002 0.044 0.004 0.037 0.000 0.012 0.000 0.010
$\alpha_{3}$ 0.001 0.309 0.012 0.266 0.005 0.091 0.000 0.073
$\gamma_{0}$0.616 3.317 -0.221 2.191 0.275 1.076 0.421 0.972
$\gamma_{1}$ 0.089 0.399 0.038 0.252 -0.015 0.107 -0.027 0.089
$\gamma_{2}$ 0.044 0.265 0.022 0.170 -0.007 0.074 -0.011 0.061
$\gamma_{3}$ 0.046 0.273 0.019 0.178 -0.008 0.078 -0.012 0.059
$\theta$0.064 0.390 -0.079 0.339 -0.119 0.260 -0.145 0.261
Runs converged (%) 99.900 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
@@
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$2.358 11.286 -1.457 8.696 0.231 4.747 0.450 4.028
$\alpha_{1}$0.360 1.770 -0.227 1.313 0.040 0.744 0.095 0.711
$\alpha_{2}$0.021 0.138 -0.012 0.142 0.002 0.063 0.006 0.065
$\alpha_{3}$0.175 0.967 -0.138 0.867 0.006 0.434 0.031 0.347
$\gamma_{0}$ 9.868 21.289 10.490 13.359 13.151 13.453 13.177 13.384
$\gamma_{1}$0.015 2.542 -0.130 1.042 -0.202 0.400 -0.186 0.339
$\gamma_{2}$0.057 1.656 -0.083 0.735 -0.090 0.274 -0.099 0.224
$\gamma_{3}$0.020 2.115 -0.065 0.719 -0.097 0.279 -0.092 0.232
$\theta$7e+05 5e+06 -3e+05 2e+06 -1e+05 4e+05 -9e+04 1e+05
Runs converged (%) 99.600 99.800 98.300 98.900
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$2.008 11.292 -1.149 8.680 0.145 4.655 0.274 3.838
$\alpha_{1}$0.308 1.773 -0.181 1.313 0.020 0.713 0.041 0.590
$\alpha_{2}$0.016 0.136 -0.007 0.139 0.003 0.057 0.004 0.043
$\alpha_{3}$0.147 0.985 -0.113 0.875 -0.003 0.429 0.017 0.333
$\gamma_{0}$ 0.661 11.910 -1.090 13.475 -0.665 4.501 -0.404 3.245
$\gamma_{1}$0.044 1.600 0.172 1.770 0.098 0.556 0.066 0.391
$\gamma_{2}$0.056 1.086 0.088 1.172 0.054 0.346 0.024 0.252
$\gamma_{3}$0.020 1.205 0.108 1.225 0.049 0.365 0.031 0.267
$\theta$0.229 0.475 -0.163 0.464 -0.090 0.410 -0.108 0.394
Runs converged (%) 99.200 99.200 99.700 100.000
Sample size ($T$) 100 200 1000 1500

### v2

@@ Acknowledgments
We thank Jeremy Fox, Yelda Gungor, and Isabelle Perrigne for their valuable comments. We were supported by JST ERATO Grant Number JPMJER2301 and Otani was supported by JSPS Grant-in-Aid (KAKENHI) for Young Researcher JSPS 24K22604 and 25K16620.
@@ Appendix A Online appendix
A.1 Summary of Goldman and Uzawa (1964)
goldmanNote1964 investigate separability concepts in demand analysis. Let $N$ be a set of variables, $N=\{1,\ldots,n\}$. Consider a partition of $N$ into $S$ parts, $\{N^{1},\ldots,N^{S}\}$ such that $N=\bigcup_{s=1}^{S}N^{s}$ and $N^{s}\cap N^{t}=\emptyset$.
A function is weakly separable with respect to a partition if
$\frac{\partial}{\partial x_{k}}\left(\frac{\partial f(x)/\partial x_{i}}{\partial f(x)/\partial x_{j}}\right)=0,\quad i,j\in N^{s},k\notin N^{s}.$ (15)
This implies that, for example, the marginal rate of substitution between $i$ and $j$ in the same partition is independent of the quantities of commodities outside $N^{s}$.
In Theorem 2, goldmanNote1964 characterizes the function that satisfies the weak separability.
Theorem 2 (Theorem 2 in goldmanNote1964).
A function $f(x)$ is weakly separable with respect to a partition $\{N^{1},...,N^{s}\}$ if, and only if, $f(x)$ is of the form:
$f(X)=\Phi(f^{1}(x^{(1)}),\ldots,f^{s}(x^{(s)}))$ (16)
where $\Phi(f^{1},\ldots,f^{s})$ is a function of $S$ variables and, for each $s$, $f^{s}(x^{(s)})$ is a function subvector $x^{(s)}$ alone.
Let $\{Q,X_{1}^{d},\ldots,X_{K}^{d}\}$ be the set of dependent variables and divide the variables into two groups $(Q,X^{d})$. Then, as the inverse demand function $P(Q,X^{d})=f(Q,r(X^{d}))$ in Theorem 1 has the form in the above theorem, we can say that the conduct parameter is not identified when the inverse demand function is separable.
In contrast, we can see that the linear demand function in bresnahan1982oligopoly and the log-linear demand function in this paper are not separable. For the linear case,
$\frac{\partial P(Q,X^{d})/\partial Z^{R}}{\partial P(Q,X^{d})/\partial Y}=\frac{\alpha_{2}Q}{\alpha_{3}}\Longrightarrow\frac{\partial}{\partial Q}\left(\frac{\partial P(Q,X^{d})/\partial Z^{R}}{\partial P(Q,X^{d})/\partial Y}\right)=\frac{\alpha_{2}}{\alpha_{3}}\neq 0.$ (17)
The last part, $\frac{\alpha_{2}}{\alpha_{3}}\neq 0$, is concluded from the identification condition in matsumura2023resolving. For the log-linear case, we have
$\frac{\partial P(Q,X^{d})/\partial Z^{R}}{\partial P(Q,X^{d})/\partial Y}=\frac{\alpha_{2}Y\log Q}{\alpha_{3}}\Longrightarrow\frac{\partial}{\partial Q}\left(\frac{\partial P(Q,X^{d})/\partial Z^{R}}{\partial P(Q,X^{d})/\partial Y}\right)=\frac{\alpha_{2}Y}{\alpha_{3}Q}\neq 0.$ (18)
Thus, again when $\frac{\alpha_{2}}{\alpha_{3}}$ is nonzero, which emphasizes the role of the demand shifters, the inverse demand function is not separable. Note that due to the log specification, $Y$ and $Q$ are positive.
A.2 Existence and uniqueness of equilibrium prices
@@
An important note is that $P_{t}=0$ satisfies $\Delta(P_{t})=0$. Thus $P_{t}^{*}=0$ holds. However, our interest is a unique positive equilibrium price, so we seek conditions in which there is a unique positive price that satisfies $\Delta(P_{t})=0$.
When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, (I) is always negative on $P_{t}>0$. In contrast, (II) is non-negative regardless of the sign of $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ on $P_{t}>0$. Therefore, $\Delta(P_{t})$ is always negative on $P_{t}>0$, which implies that there is no positive equilibrium price.
@@
Therefore, the shape of (II) on $P_{t}>0$ is changed based on the value of $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ and the number that (I) and (II) cross on $P_{t}>0$ is also determined.
The case $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})<0$ is illustrated in Figure 1(a). In this case, (II) is a monotone decreasing convex function because the first and second derivatives are both negative. As (I) is monotone increasing in $P_{t}>0$, (I) and (II) cross only once on $P_{t}>0$.
The case $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\in[0,1)$ is illustrated in Figure 1(b). In this case, (II) becomes a monotone increasing concave function passing through the origin because the first derivative is positive and the second derivative is negative. As the value of the first derivative is infinity at $P_{t}=0$ since $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})-1<0$, (II) should be greater than (I) around $P_{t}=0$. Then, as $P_{t}$ becomes large starting from zero, the difference in both terms becomes small, and both terms cross only once on $P_{t}>0$ eventually.
When $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, (20) becomes
@@
If $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})=\exp(\Xi)$, the above equation holds for any $P_{t}$. Thus, there are infinitely many equilibrium prices on $P_{t}>0$, which is illustrated in Figure 1(c). When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq\exp(\Xi)$, there is no equilibrium price on $P_{t}>0$ because the right-hand side of the equation is always non-zero on $P_{t}>0$, which is illustrated in Figure 1(d).
The case $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})>1$ is illustrated in Figure 1(e). In this case, (II) becomes a monotone increasing convex function passing through the origin because the first and second derivatives are positive. Therefore, (I) and (II) cross only once on $P_{t}>0$.
@@
$P_{t}$$\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$$1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})]P_{t}$
@@
Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>0$. The thick line represents the first term and the dotted line represents the second term in (20).
@@
$0$ $=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})]P_{t}-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$
$0$ $=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$
@@
Figure 2 illustrates how the demand and supply equations cross under the conditions. When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the supply equation (5) is ill-defined because the inside of the log function becomes negative. Thus, there should not be any equilibrium. Hereafter, assume that $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>0$. When $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand equation and supply equation coincides. Thus there are infinitely many equilibria in the model. When $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand and supply equations have a same slope and different intercepts, which means that both equations become parallel. Thus, there is no equilibrium. When $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, the demand and supply equations have different slopes, and hence we can find a unique equilibrium.
@@
$Q_{t}$$P_{t}$SSDD
(a) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
$Q_{t}$$P_{t}$SSDD
(b) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
$Q_{t}$$P_{t}$SSDD
@@
Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>0$. Case (a) and (b) also allow both the demand and supply equations to be downward sloping.
A.3 Simulation and estimation procedure
@@
We compute the equilibrium quantity $Q_{t}$ for the log-linear model by (24). We then compute the equilibrium price $P_{t}$ by substituting $Q_{t}$ and other variables into the demand function (2). We generate 1000 data sets of 100, 200, 1000, 1500 markets. We jointly estimate the demand and supply parameters by the simultaneous equation model (wooldridge2010econometric) from the true values. We use state-of-the-art constrained optimization solvers, i.e., Ipopt.jl which implements an interior point line search filter method that aims to find a local solution of nonlinear programming problems.
@@
The ad-hoc estimation for Table 2 is formulated as follows. Let $\Phi=(\xi,\{MC_{t}\}_{t=1,\ldots,T})$. Consider a constrained minimization problem such that
$\Phi^{*}=\arg\min_{\Phi}\ g(\Phi)^{\top}Wg(\Phi),$ (25)
subject to
$(\Phi)=\left[\begin{array}[]{l}\frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{d}_{t}(\Phi)Z_{t}^{d}\\ \frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{c}_{t}(\Phi)Z_{t}^{c}\end{array}\right],$ (29)
${\varepsilon}_{t}^{d}(\Phi)=\log P_{t}-\alpha_{0}+(\alpha_{1}+\alpha_{2}Z^{R}_{t})\log Q_{t}-\alpha_{3}\log Y_{t},$ (30)
${\varepsilon}_{t}^{c}(\Phi)=\log MC_{t}-\gamma_{0}-\gamma_{1}\log Q_{t}-\gamma_{2}\log W_{t}-\gamma_{3}\log R_{t},$ (31)
$P_{t}=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+MC_{t},$ (32)
$MC_{t}\geq 0,$ (33)
where the matrix $W$ is the same as the original N2SLS.
@@
A.4 Why does the estimation without equilibrium conditions converge to the extremely low conduct parameters?
In the literature, it is often observed that estimation without imposing equilibrium conditions tends to converge to negative values of the conduct parameter. We confirm this pattern in our Monte Carlo simulations, particularly in Panel (a) of Table 1. To understand this phenomenon, consider the supply equation (5), which includes the following term:
$-\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}{t}))+\gamma{0}.$ (34)
If the model allows for $\theta<0$, the term inside the logarithm can become arbitrarily large in magnitude by choosing a sufficiently negative $\theta$ when $\alpha_{1}+\alpha_{2}Z^{R}_{t}>0$:
$\lim_{\theta\rightarrow-\infty}\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))\rightarrow\infty.$ (35)
However, since $\theta$ appears inside a logarithmic function, changes in $\theta$—even when large—have a relatively small impact on the GMM objective function. Moreover, the effect of changes in $\theta$ can be offset by adjusting $\gamma_{0}$, with only modest changes in its value. This implies that the GMM objective function can be very flat in the neighborhood of the true values of $\theta$ and $\gamma_{0}$.
We verify this implication by examining the shape of the GMM objective function. Figure 3 plots the objective function on the $\theta$$\gamma_{0}$ plane, fixing all other parameters at their true values and using one simulated dataset. The same pattern is observed across other simulation datasets as well. As the figure clearly shows, the GMM objective function is nearly flat around the true value of $\gamma_{0}$, and it allows for a wide range of $\theta$ values, especially in the negative direction, without significantly affecting the objective function value.
Figure 3: The contour map of the GMM objective function on $\theta-\gamma_{0}$ plane.
Note: We pick one simulation data where $T=100$ and $\sigma=1$. The GMM objective function is flat around the true values of the intercept of the log marginal cost and the conduct parameter.
A.5 Additional experiments
@@
Additional results for different $\sigma$ are shown in Tables 4 and 5. As a summary, the main findings in the main text are robust.
Table 6 shows the results of the estimation with ado hoc method with different standard deviations of the error terms. As in Table 2, the run convergence rate becomes 100% and the accuracy of the conduct parameter estimation is improved. However, again, the results with ad hoc improvement do not dominate the results without the improvement for all parameters.
@@
In Tables 7 and 8, we illustrate that N2SLS with Constraints (12) works for the linear model as in matsumura2023resolving. This means that incorporating the conditions is innocuous for the linear model.
Table 9 demonstrates that relying solely on Constraints (13) and (14) within the N2SLS framework leads to severe bias, implying that only Constraints (13) and (14) cannot resolve the problem. Table 10 demonstrates that relying solely on Constraints (12) within the N2SLS framework leads to severe bias—especially in small samples—for parameters other than the conduct parameter, as well as poor convergence. While the conduct parameter itself appears stable, this is mechanically due to the imposed domain constraint, which prevents extreme estimates by construction. These results underscore that Constraints (12) alone are insufficient and highlight the critical importance of incorporating the equilibrium conditions (13) and (14). In large samples, however, the problem is substantially alleviated, as the parameter search becomes less problematic.
@@
(a)N2SLS without Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ 0.119 4.055 0.346 3.180 0.344 1.497 0.379 1.582 $\alpha_{1}$ 0.019 0.620 0.054 0.486 0.076 0.429 0.099 0.576 $\alpha_{2}$ 0.000 0.050 0.003 0.039 0.002 0.029 0.003 0.036 $\alpha_{3}$ 0.009 0.336 0.025 0.268 0.028 0.129 0.031 0.129 $\gamma_{0}$ 10.104 11.066 9.397 10.288 13.240 13.280 13.314 13.346 $\gamma_{1}$ -0.146 0.340 -0.166 0.264 -0.180 0.198 -0.181 0.192 $\gamma_{2}$ -0.072 0.229 -0.079 0.165 -0.087 0.108 -0.086 0.099 $\gamma_{3}$ -0.069 0.230 -0.082 0.171 -0.088 0.110 -0.087 0.100 $\theta$ -5e+04 1e+05 -3e+04 6e+04 -8e+04 9e+04 -8e+04 9e+04 Runs converged (%) 100.000 100.000 98.100 98.700 Sample size ($T$) 100 200 1000 1500
(b) N2SLS with Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.006 3.556 0.175 3.158 0.059 0.977 -0.017 0.763 $\alpha_{1}$ -0.002 0.545 0.026 0.484 0.009 0.150 -0.003 0.117 $\alpha_{2}$ 0.002 0.044 0.004 0.037 0.000 0.012 0.000 0.010 $\alpha_{3}$ 0.001 0.309 0.012 0.266 0.005 0.091 0.000 0.073 $\gamma_{0}$ -0.616 3.317 -0.221 2.191 0.275 1.076 0.421 0.972 $\gamma_{1}$ 0.089 0.399 0.038 0.252 -0.015 0.107 -0.027 0.089 $\gamma_{2}$ 0.044 0.265 0.022 0.170 -0.007 0.074 -0.011 0.061 $\gamma_{3}$ 0.046 0.273 0.019 0.178 -0.008 0.078 -0.012 0.059 $\theta$ -0.064 0.390 -0.079 0.339 -0.119 0.260 -0.145 0.261 Runs converged (%) 99.900 100.000 100.000 100.000 Sample size ($T$) 100 200 1000 1500
@@
(a) N2SLS without Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -2.358 11.286 -1.457 8.696 0.231 4.747 0.450 4.028 $\alpha_{1}$ -0.360 1.770 -0.227 1.313 0.040 0.744 0.095 0.711 $\alpha_{2}$ -0.021 0.138 -0.012 0.142 0.002 0.063 0.006 0.065 $\alpha_{3}$ -0.175 0.967 -0.138 0.867 0.006 0.434 0.031 0.347 $\gamma_{0}$ 9.868 21.289 10.490 13.359 13.151 13.453 13.177 13.384 $\gamma_{1}$ -0.015 2.542 -0.130 1.042 -0.202 0.400 -0.186 0.339 $\gamma_{2}$ -0.057 1.656 -0.083 0.735 -0.090 0.274 -0.099 0.224 $\gamma_{3}$ -0.020 2.115 -0.065 0.719 -0.097 0.279 -0.092 0.232 $\theta$ -7e+05 5e+06 -3e+05 2e+06 -1e+05 4e+05 -9e+04 1e+05 Runs converged (%) 99.600 99.800 98.300 98.900 Sample size ($T$) 100 200 1000 1500
(b) N2SLS with Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -2.008 11.292 -1.149 8.680 0.145 4.655 0.274 3.838 $\alpha_{1}$ -0.308 1.773 -0.181 1.313 0.020 0.713 0.041 0.590 $\alpha_{2}$ -0.016 0.136 -0.007 0.139 0.003 0.057 0.004 0.043 $\alpha_{3}$ -0.147 0.985 -0.113 0.875 -0.003 0.429 0.017 0.333 $\gamma_{0}$ 0.661 11.910 -1.090 13.475 -0.665 4.501 -0.404 3.245 $\gamma_{1}$ -0.044 1.600 0.172 1.770 0.098 0.556 0.066 0.391 $\gamma_{2}$ -0.056 1.086 0.088 1.172 0.054 0.346 0.024 0.252 $\gamma_{3}$ -0.020 1.205 0.108 1.225 0.049 0.365 0.031 0.267 $\theta$ -0.229 0.475 -0.163 0.464 -0.090 0.410 -0.108 0.394 Runs converged (%) 99.200 99.200 99.700 100.000 Sample size ($T$) 100 200 1000 1500